OpenStudy (anonymous):

use the given values to write an equation that relates x and y. then find y when x=2.

OpenStudy (anonymous):

x=8,y=-1

OpenStudy (anonymous):

Form of the equation is y=x-9 so at x=2 ,y=-7

OpenStudy (anonymous):

Can you answer this? and do you have a better answer for my first question?

OpenStudy (anonymous):

number of chips per wafer is area is 79 is=26000/79